# On the optimality of coin-betting for mean estimation

### v2

@@ 5 Optimality beyond single-round e-variables
Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{[0,1]}$, which cannot capture dependencies across multiple rounds. In such setting, the dependency structure among observations is fixed and assumed a priori, rather than being subject to testing. Yet, e-variable-based testing naturally extends to more general hypotheses that span multiple rounds and can account for sequential or dependent data structures. In the second part of the paper, we turn our attention to this richer framework. Remarkably, a slight adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$. We will first show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes. Conversely, we will show that when testing the i.i.d. assumption, the coin-betting approach no longer yields the optimal e-class.
For technical reasons (primarily related to measurability, see discussion in Section 7), for the rest of this work we restrict ourselves to the case of a random variable taking values in a finite (known) set. Therefore, henceforth we let $\mathcal{X}\subseteq\mathbb{R}$ denote a finite set. For simplicity, we assume that $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$, although all our results extend readily to any case where the minimum and maximum values are finite. We note that since $\mathcal{X}$ has finite cardinality, every function $\mathcal{X}^{t}\to\mathbb{R}$ is automatically Borel measurable.
5.1 Optimal e-class for multi-round e-variables
@@
where $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$. Note that $\mathcal{H}_{\mu}^{T}$ is a hypothesis on $\mathcal{X}^{T}$. As such, its set of e-variables will consist of Borel functions from $\mathcal{X}^{T}$ to $\mathbb{R}$. We denote the set of all e-values relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
In a coin-betting game, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$. It is straightforward to check that under the hypothesis $\mathcal{H}_{\mu}^{T}$, the expectation of $M_{T}$ is always exactly one ($M_{T}$ is a martingale), which shows that $M_{T}$ is an e-variable. Next, we show that the e-class of e-variables in this form coincides is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
@@
$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
@@
The proof of Theorem 2 shares many similarity with that of Theorem 1. In particular, we use the same technique to establish the maximality of e-variables of the form $E_{\lambda^{T}}$. A geometric argument analogous to the one illustrated in Figure 1 is then employed, as part of an induction recursion, to prove that the e-class of interest is majorising. However, the detailed proof being somewhat lengthy, we defer it to the appendix.
@@
From what we have shown so far, in a $T$ round game coin-betting yields all and only the maximal e-variables, when testing the hypothesis that the conditional mean is some fixed $\mu\in(0,1)$. Yet, this is not any more the case if we consider a more restrictive hypotheses, stating that the observations are i.i.d. and with mean $\mu$. More concretely, let
@@
To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, we consider the “highly symmetrical” case of $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$. We show at the end appendix that the e-variables in this setting are the functions $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ that satisfy
@@
where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$. Clearly, the optimal e-class here consists of those e-variables satisfying the constraint on $\xi_{1}$ with equality. We note that all the e-variables in $\mathcal{E}_{1/2}^{T}$ satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$. As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$, but there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$. The function that equals $4$ on $x^{2}=(1,1/2)$, and 00 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
@@
We leave as a non-trivial open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
6 E-processes for testing the fixed conditional mean
@@
A key advantage of many testing procedures involving e-variables is anytime validity. One way to achieve this is through the round-wise testing-by-betting approach, which gives rise to a super-martingale (cf. Lemma 1). However, as we have already pointed out, this is not the most general approach. Rather than only focusing on super-martingales, one can instead use any non-negative process whose expectation is upper bounded by one, regardless the stopping time. These are known as e-processes, a generalization of super-martingales, which can enable powerful tests even in scenarios where martingale-based methods lack power (Ramdas et al., 2022b).
@@
As in the previous section, here $\mathcal{X}$ denotes a finite subset of $[0,1]$, with $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$. We denote as $\mathcal{X}^{\infty}$ the space of sequences $(x_{t})_{t\geq 1}\subseteq\mathcal{X}$. For convenience, we will sometimes write $x^{\infty}$ to mean $(x_{t})_{t\geq 1}$. We endow the set $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$. For a fixed $\mu\in(0,1)$, in this section we will focus on the hypothesis class
$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,.$
@@
This time, we will not aim to study the e-variables for $\mathcal{H}^{\infty}_{\mu}$, as these would be functions that take a whole sequence as argument. Conversely, we are interested in tools that allow us to stop the test if we think enough data-points have been observed to decide whether or not the hypothesis has to be rejected. To make this rigorous we first need some definitions. First, we call a finite stopping time a function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, whenever for some $x^{\infty}$ and $y^{\infty}$ in $\mathcal{X}^{\infty}$ and some $t\in\mathbb{N}$ we have $\tau(x^{\infty})=\tau(y^{\infty})=t$, then $x^{t}=y^{t}$. [^8] It is straightforward to verify that in our setting this definition is equivalent to the perhaps more classical notion of finite e-process as a $\tau$ such that the sets $\{\tau\leq t\}$ are $\mathcal{F}_{t}$ measurable. We say that $\tau$ is bounded by $T\geq 0$ if its image is contained in $[0:T]$. We let $\mathcal{T}_{T}$ denote the set of all stopping times bounded by $T$.
Definition 6.
@@
An e-process (for $\mathcal{H}_{\mu}^{\infty}$) is a sequence $(E_{t})_{t\geq 0}$, of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that, for every $P\in\mathcal{H}^{\infty}_{\mu}$, any $T\geq 0$ and any $\tau\in\mathcal{T}_{T}$,
$\mathbb{E}_{P}[E_{\tau}]\leq 1\,.$
@@
We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all the e-processes for $\mathcal{H}_{\mu}^{\infty}$. We call e-process class any subset of $\mathcal{E}_{\mu}^{\infty}$. Similarly to what we have done for the e-variables, we say that the e-process $E$ majorises the e-process $E^{\prime}$ (we write $E\succeq E^{\prime}$) if for every $t\geq 0$ we have $E_{t}\succeq E^{\prime}_{t}$. We say that an e-process class $\mathcal{E}$ is majorising if for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$. We say that $\mathcal{E}$ is optimal if it is majorising and is contained in every other majorising e-process class. We call an e-process $E$ maximal if it is not strictly majorised by any other e-process. With the same reasoning as in Lemma 3, one can show that the optimal e-process class exists if, and only if, the class of all the maximal e-processes is majorising. In such case the optimal class is unique and coincides precisely with all and only the majorising e-processes.
We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, such that $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$. For any $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$, we define $E^{\lambda^{\infty}}=(E^{\lambda^{\infty}}_{T})_{T\geq 0}$ via
@@
for $T\geq 1$, and $E_{0}=1$. These are the e-processes associated to the wealth of a coin-betting player. As $E^{\lambda^{\infty}}$ defines a martingale under $\mathcal{H}_{\mu}^{\infty}$, it is also an e-process for such hypothesis. We define the coin-betting e-process class $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}$.
The next result shows that for $\mathcal{H}_{\mu}^{\infty}$ the coin betting approach is optimal also at the level of e-processes. The proof (reported in the appendix) always follows the ideas introduced in the proof of Theorem 1.
Theorem 3.
For any $\mu\in(0,1)$, $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
@@ 7 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$. Theorem 2 and Theorem 3 extended this optimality result to testing the fixed conditional mean of a sequence of bounded observations. To formalise these claims, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this paper. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
@@
The concept of maximality for e-variables, as introduced in this paper, is essentially equivalent to the classical statistical notion of admissibility. An e-variable is considered maximal if no other e-variable strictly dominates it. Likewise, the idea of a majorising e-class corresponds to the classical concept of a complete class, with the optimal e-class representing the minimal complete class of e-variables. For a detailed discussion of these ideas in traditional statistics we refer to Lehmann and Romano (2022). In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the classical partial order of dominance among functions.
@@
The concept of admissibility in the context of e-variables and e-processes was previously introduced by Ramdas et al. (2022a), whose definition is closely aligned with our notion of maximality. The key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance. [^9] The two definitions becomes of course equivalent when $\mathcal{X}$ is finite, but they are not in the case of $\mathcal{X}=[0,1]$, which is the setting of our analysis on single-round e-variables. This stricter requirement is motivated by the will of developing a theory that remains valid even in the absence of a known alternative, where any value in $\mathcal{X}$ might occur at the next observation. Although Ramdas et al. (2022a) provide some necessary and sufficient conditions for admissibility, we emphasize that their results are insufficient to establish the optimality of the coin-betting e-class. Specifically, their approach assumes the existence of a reference measure for the hypothesis, [^10] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a condition not satisfied by $\mathcal{H}_{\mu}$. More precisely, their results can show that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class. This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables (or equivalently, wealth processes), rather than evaluating them in isolation. It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees nothing is lost by relying on the coin-betting approach. This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually. Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
@@
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$. However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$. Although one might argue that for each of the maximal e-variable there is at least one measure that bring the expectation to one (Grünwald, 2024 calls such property sharpness), sharpness alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$ is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).

### v4

@@ 6 Optimality beyond single-round e-variables
Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{\mathcal{X}}$, which cannot capture depedences across multiple rounds. In such setting, the dependence structure among observations was fixed and assumed a priori, rather than being subject to testing. Yet, e-variable-based testing naturally extends to more general hypotheses that span multiple rounds and can account for sequential or dependent data structures. In the second part of the paper, we turn our attention to this richer framework. Remarkably, an adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$. We show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes. Conversely, we will show that when testing the i.i.d. assumption, the coin-betting approach no longer yields the optimal e-class.
6.1 Optimality with multi-round e-variables
@@
where $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$. Note that $\mathcal{H}_{\mu}^{T}$ is a hypothesis on $\mathcal{X}^{T}$. As such, its set of e-variables will consist of Borel functions from $\mathcal{X}^{T}$ to $\mathbb{R}$. We denote the set of all e-variablues relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
Let us consider a test where all the $T$ observations $x^{T}$ are seen at once, where a single e-variable $E\in\mathcal{E}_{\mu}^{T}$ needs to be selected. If $E(x^{T})\geq 1/\delta$, the null $\mathcal{H}^{T}_{\mu}$ is rejected. We remark that this setting departs from the sequential testing-by-betting games discussed in the previous sections, as now the data set size $T$ is fixed in advance and the player makes a single decision before seeing the entire data set. In a way, this is as if we where playing a single round in the game of Definition 1, with $\mathcal{Z}=\mathcal{X}^{T}$. [^8] Of course, this can also be extended in a sequential game, where each round a new block of $T$ data points is observed. Despite this difference, the connection to the $T$-round coin-betting game of Definition 2 remains strong. Indeed, in the coin-betting setup that we have considered earlier, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$. It is straightforward to check that under the hypothesis $\mathcal{H}_{\mu}^{T}$, the expectation of $M_{T}$ is always exactly one ($M_{T}$ is a martingale), which shows that $M_{T}$ is an e-variable. Next, we show that the e-class of e-variables in this form coincides with the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
@@
$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a Borel function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
@@
The proof of Theorem 2 shares many similarities with that of Theorem 1. In particular, we use the same technique to establish the maximality of e-variables of the form $E_{\lambda^{T}}$. A geometric argument analogous to the one illustrated in Figure 1 can then be employed, as part of an induction recursion, to prove that the e-class of interest is majorising. However, the detailed proof is somewhat lengthy and involves some technical subtleties when dealing with measurability. We hence defer it to A.3.
@@
From what we have established so far, coin-betting yields all and only the maximal e-variables when testing the $T$-round hypothesis $\mathcal{H}_{\mu}^{T}$ that the conditional mean is some fixed $\mu\in(0,1)$. Yet, this is not any more the case if we consider a more restrictive hypothesis, stating that the observations are i.i.d. and with mean $\mu$. More concretely, let us define
@@
To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, let us consider the “highly symmetrical” case $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$. We show in A.2 that the e-variables in this setting are the $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ satisfying
@@
where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$. Clearly, the optimal e-class here consists of those e-variables that satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$, or $\xi_{1}=1+\sqrt{(1-\xi_{0})(1-\xi_{2})}$ with $\xi_{0}<1$ and $\xi_{2}<1$. We note that all the e-variables in $\mathcal{E}_{1/2}^{T}$ satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$. As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$. Yet, there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$. The function that equals 444 on $x^{2}=(1,1/2)$, and 0 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
@@
We leave as a (non-trivial) open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
6.3 Optimal classes of e-processes
@@
A key advantage of many testing procedures involving e-variables is allowing early stopping. One way to achieve this is through the round-wise testing-by-betting approach that we presented in Section 2, which gives rise to a super-martingale (cf. Proposition 1). However, as we have already pointed out, this is not the most general approach. Rather than only focusing on super-martingales, one can in principle use any non-negative process whose expectation is upper bounded by one, regardless the stopping time. These are known as e-processes, a generalization of super-martingales, which can enable powerful tests even in scenarios where martingale-based methods lack power (Ramdas et al., 2022b).
@@
We denote as $\mathcal{X}^{\infty}$ the space of sequences $(x_{t})_{t\geq 1}\subseteq\mathcal{X}$. For convenience, we will often write $x^{\infty}$ for $(x_{t})_{t\geq 1}$. We endow $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$. For a fixed $\mu\in(0,1)$, in this section we focus on the hypothesis class
$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,,$
@@
where again we use the convention $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$. This time, we will not aim to study the e-variables for $\mathcal{H}^{\infty}_{\mu}$, as these would be functions that take a whole sequence as argument. Conversely, we are interested in tools that allow us to stop the test if we think enough data-points have been observed to decide whether or not the hypothesis has to be rejected. To make this rigorous we first need some definitions. First, we define a finite stopping time as a measurable [^13] Here we endow $\mathbb{N}$ with the discrete sigma-field. function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, for any $t\geq 0$ the set $\{\tau=t\}$ is measurable with respect to the sigma-field $\mathcal{F}_{t}$, generated by the projection on the first $t$ components of the sequences in $\mathcal{X}^{\infty}$. We let $\mathcal{T}$ denote the set of all finite stopping times.
Definition 6
@@
Let $E=(E_{t})_{t\geq 0}$ be a sequence of non-negative Borel functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$. We say that $E$ is an e-process (for $\mathcal{H}_{\mu}^{\infty}$) when, for any $P\in\mathcal{H}^{\infty}_{\mu}$ and any $\tau\in\mathcal{T}$, $\mathbb{E}_{P}[E_{\tau}]\leq 1$.
Sequential testing using e-processes proceeds as follows (see Ramdas and Wang 2024 for more details). Before observing any data, an e-process $E=(E_{t})_{t\geq 0}$ for $\mathcal{H}_{\mu}^{\infty}$ is fixed. [^14] Note that although the e-process is fixed before observing any data, each $E_{t}$ is a function of the past observations $x^{t-1}$, allowing the test to adapt to the data as they become available. Then, at each round $t$, the player’s wealth is defined as $\log E_{t}(x^{t})$. $\mathcal{H}_{\mu}^{\infty}$ is rejected if this wealth ever exceeds $\log 1/\delta$. The definition of an e-process ensures that this yields a sequential test with Type I error controlled at level $\delta$. We remark that this testing procedure extends the testing-by-betting framework of Definition 1 to the more expressive setting where the null hypothesis takes into account the entire dependence structure of the data sequence.
@@
We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all e-processes for $\mathcal{H}_{\mu}^{\infty}$. An e-process class is any subset of $\mathcal{E}_{\mu}^{\infty}$. Similarly to what we have done for the e-variables, we say that the e-process $E$ majorises the e-process $E^{\prime}$ (we write $E\succeq E^{\prime}$) if for every $t\geq 0$ we have $E_{t}\succeq E^{\prime}_{t}$. If $E\succeq E^{\prime}$ and there is a $t\geq 0$ such that $E_{t}\succ E^{\prime}_{t}$, then $E$ strictly majorises $E^{\prime}$, and we write $E\succ E^{\prime}$. We call an e-process maximal if it is not strictly majorised by any other e-process. We say that an e-process class $\mathcal{E}$ is majorising if, for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$. We say that $\mathcal{E}$ is optimal if it is majorising and all its elements are maximal. We remark that if an optimal majorising e-process class exists, then it is unique, it consists with the set of all the maximal e-processes, and it is included in every other majorising e-process class.
We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, with $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$ Borel for all $t\geq 1$. For any $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$, we define $E^{\lambda^{\infty}}=(E^{\lambda^{\infty}}_{T})_{T\geq 0}$ via
@@
for $T\geq 1$, and $E_{0}=1$. These are the sequences associated with the wealth of a coin-betting player. As $E^{\lambda^{\infty}}$ defines a martingale under $\mathcal{H}_{\mu}^{\infty}$, it is also an e-process for such hypothesis. We hence define the coin-betting e-process class
$\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}\,.$
The next result shows that the coin-betting approach is optimal at the level of e-processes for $\mathcal{H}_{\mu}^{\infty}$. The proof (see A.4) follows the ideas introduced in the proofs of Theorems 1 and 2.
Theorem 3
For any $\mu\in(0,1)$ the coin-betting e-process class is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
@@ 7 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to test the mean and build confidence sequences given a sequence of independent draws from an unknown supported on [0,1][0,1][ 0 , 1 ]. Theorems 2 and 3 extend this optimality result when testing the hypothesis of a fixed conditional mean. To formalise these claims, we introduced the notions of majorising and optimal e-classes, which may be of independent interest in the context of sequential testing, beyond the scope of this paper. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024). Notably, log-optimality is defined in terms of an alternative hypothesis, against which the null is tested. In such setting there can be indeed a single (up to null sets under the alternative hypothesis) best e-variable. However, here we adopt a different perspective, where no alternative is defined. It is not hard to see that if the optimal e-class exists, whenever one considers an alternative that allows to define a log-optimal e-variable, a “version” [^15] As the log-optimal e-variable is defined up to null sets under the alternative, there might be elements of this family of e-variables that are not maximal. However, there is always a maximal e-variable among them if the optimal e-class exists. of this e-variable must lie in the optimal e-class. Interestingly, in the case of the coin-betting e-class, for each $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ one can find an alternative hypothesis such that $E$ is log-optimal. [^16] Specifically, Theorem 1 in Grünwald et al. (2024) implies that, for $\lambda\in I_{\mu}$, $E_{\lambda}$ is log-optimal for the point alternative $\{Q\}$, with $Q=(1-\mu)E(0)\delta_{0}+\mu E(1)\delta_{1}$, since $E_{\lambda}=\mathrm{d}Q/\mathrm{d}P$ on the support of $Q$, where $P=(1-\mu)\delta_{0}+\mu\delta_{1}$ is in $\mathcal{H}_{\mu}$.
@@
The concept of maximality for e-variables, as introduced in this paper, is essentially equivalent to the classical statistical notion of admissibility. An e-variable is considered maximal if no other e-variable strictly dominates it. Likewise, the idea of a majorising e-class corresponds to the notion of a complete class, with the optimal e-class representing the minimal complete class of e-variables. In this sense, the problem of identifying the optimal e-class can be seen as an instance of characterising the minimal complete class of tests, a question largely studied in classical statistics. For a detailed discussion of these ideas in traditional statistics we refer to Lehmann and Romano (2022). In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the standard dominance partial ordering among functions.
@@
The concept of admissibility in the context of e-variables and e-processes was previously introduced by Ramdas et al. (2022a), whose definition is closely aligned with our notion of maximality. However, a key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance. This stricter requirement is motivated by the will of developing a theory that remains valid even in the absence of a known alternative, where any value in $\mathcal{X}$ might occur at the next observation. Although Ramdas et al. (2022a) provide some necessary and sufficient conditions for admissibility, we emphasise that their results are insufficient to establish the optimality of the coin-betting e-class. Specifically, their necessary conditions assume the existence of a reference measure for the considered hypothesis, [^17] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a requirement not satisfied by $\mathcal{H}_{\mu}$. More precisely, their findings imply that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class. This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables, rather than evaluating them in isolation. It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees that nothing is lost by relying on the coin-betting approach.
@@
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noting that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals 111 under every $P\in\mathcal{H}_{\mu}$. However, in general this property (often called exactness) merely implies that all the elements in the e-class are maximal, and not its optimality. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$ (see Clerico, 2024). Although one might argue that for each maximal e-variable there is at least one measure that brings the expectation to 111 (Grünwald 2024 calls such property sharpness), this feature alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$, the function equal to 111 on $\mu$ and 0 elsewhere, is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).